40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Mar 12 01:33:21 2020, gautam, Update, LSC, Resuming locking activities ExcessFreqNoise_LSC.pdf
    Reply  Fri Mar 13 00:32:38 2020, gautam, Update, LSC, Some progress RFonly_CARM.pngCARMTF.pdf
       Reply  Tue Mar 17 01:22:03 2020, gautam, Update, LSC, Locking updates 
          Reply  Thu Mar 19 03:33:28 2020, gautam, Update, LSC, More locking updates PRFPMI_lock.pngCARMTF.pdfDARM_AvB.pdflockLoss.png
Message ID: 15278     Entry time: Tue Mar 17 01:22:03 2020     In reply to: 15273     Reply to this: 15281
Author: gautam 
Type: Update 
Category: LSC 
Subject: Locking updates 

Summary:

No real progress tonight - I made it a bunch of times to the point where CARM was RF only, but I never got to run a measurement to determine what the DARM_B loop gain should be to make the control fully RF.

Details:

  • Touched up PMC alignment.
  • There were very few BNC cables available at the rack near SW corner of the PSL table - the short BNC cables are NOT meant to be daisy chained to make long cables to run along the arm, I removed all those.
  • Restored SR785 at LSC rack for CARM TF measurements.
  • I was able to get the CARM UGF ~5 kHz, but everytime I was trying to run a DTT swept sine to measure the ratio of DARM_B_IN1 / DARM_A_IN1, the lock was lost - not sure if this is because of the excitation injected or something else.
  • I'll probably give this another shot Wednesday eve.
ELOG V3.1.3-